### v1

@@ Abstract
LLM agents excel in settings where the environment is largely static and the relevant information fits within the model’s context window, but struggle in open-ended investigations where explanations must be derived by iteratively mining evidence from massive, heterogeneous raw data. Such settings often contain hidden dependency structure: entities interact, signals co-vary, and the relevance of a detail may only become apparent after additional discoveries elsewhere. Bounded context window further forces agents to summarize and compress intermediate findings before their significance is known, increasing the chance that crucial evidence is discarded prematurely.
ReAct-style LLM agents are particulary brittle here for two structural reasons. First, they operationalize investigations as a linear retrieve $\rightarrow$  summarize $\rightarrow$ reason loop in which each step is effectively a conditional query on the current (partial) evidence set; this makes outcomes highly sensitive to exploration order in iterative investigation and discovery tasks, introduces run-to-run non-determinism, and creates a reliability gap where Pass@k can be high while Majority@k remains much lower, indicating success by chance rather than consistent reasoning. In these settings, simply scaling inference (e.g., sampling more rollouts or longer chains) does not reliably stabilize conclusions, since candidate explanations cannot be checked or grounded autonomously. As new evidence arrives, earlier interpretations must be revised, but ReAct provides no explicit mechanism for belief bookkeeping and revision. Second, ReAct entangles semantic reasoning with controller responsibilities (tool orchestration, state tracking, and execution hygiene), so execution failures (e.g., argument errors, repetition, plan drift) leak into and degrade the reasoning process while wasting scarce context budget.
@@
1 Introduction
Large language model (LLM) agents have demonstrated strong performance on structured reasoning benchmarks—code synthesis, mathematical proof, and symbolic planning—yet remain brittle on open-ended queries where explanations must be inferred from massive heterogeneous data. This challenge arises ubiquitously across domains, including medical diagnosis, forensic investigations, and modern IT operations. In the IT domain in particular—spanning SRE, FinOps, and DevOps—cloud infrastructure, microservice architectures, CI/CD pipelines, financial platforms, and security monitoring systems continuously generate large volumes of observability data, including logs, metrics, distributed traces, configuration states, and events. When anomalies occur—service degradations, budget overruns, security incidents, or release failures—operators must navigate this ocean of data to answer a deceptively simple question: what explains the observed behavior of the system? Answering such questions is at the heart of operational diagnosis.
Formally, operational diagnosis can be cast as an abductive reasoning problem over graphs. Given a set of observed symptoms (alerts, events, failures) from a dynamic environment—which are often partial, noisy, and evolving—the goal is to infer a concise and coherent explanation that accounts for these observations in terms of the underlying system structure and dynamics.
Limits of existing agent patterns:
Current agentic approaches struggle to automate operational diagnosis. While the limitations of ReAct-style agents—such as tool invocation failures and redundant or degenerate tool calls—are well known Cemri et al. (2025), the dominant challenge in operational diagnostics lies in context gathering and sequencing under uncertainty. Reasoning is only as sound as the evidence it conditions on: if relevant context is not retrieved, it is ignored and cannot be reasoned over, regardless of model capacity. In open-ended investigations, however, the relevance of evidence is not known a priori, and ReAct agents lack principled mechanisms for ensuring systematic context acquisition. Tool calls are issued opportunistically, leading to missed low-frequency but high-impact signals or excessive retrieval of irrelevant data. A common mitigation is to summarize retrieved evidence before further reasoning, but this introduces a fundamental failure mode: summarization necessarily discards information before its relevance can be established, causing critical evidence to be lost prematurely Liu et al. (2024).
More fundamentally, operational diagnosis exhibits non-monotonic belief revision behaviors: the interpretation of an entity or signal naturally changes as new evidence is discovered elsewhere in the system, making reasoning order-dependent. As a result, the sequence in which evidence is retrieved, summarized, and presented to the LLM materially affects conclusions. In practice, ReAct agents exhibit high variance across runs due to planning non-determinism, unstable exploration orderings, and lossy intermediate summarization, producing explanations that are incomplete, irreproducible, and fragile. These failures are architectural: ReAct implicitly assumes linear, monotonic reasoning over an accumulating context, whereas operational diagnostics require structured context acquisition with explicit support for ordering and belief revision.
Towards graph-guided agentic explanations:
Building on the abductive framing above, we introduce an operational-graph abstraction for open-ended investigations under partial observability. Across IT domains such as reliability, cost management, security, and deployment, abductive explanations are naturally structured by dependency relationships among system entities. This motivates representing the environment as an operational graph and defining diagnosis as explanations over graphs: the agent reasons over a partial, evolving world model, while explanations correspond to minimal subgraphs consistent with observed evidence. Framing diagnosis in this way converts an ill-defined exploration problem into structured inference over a world model.
Crucially, diagnosis is inherently non-monotonic—beliefs about the explanatory role of an entity may change as new evidence is uncovered elsewhere in the graph. Drawing inspiration from belief propagation Pearl (1988), we model diagnosis as an iterative process in which local beliefs are assigned, revised, and propagated through the graph. This perspective naturally leads to a disaggregated (neurosymbolic) architecture in which the language model performs local reasoning over bounded neighborhoods, while a symbolic controller manages traversal order, context gathering, belief bookkeeping, and termination. This separation is essential for supporting non-monotonic belief revision: beliefs are recorded explicitly and revised as new evidence is incorporated, rather than being implicitly overwritten in a growing prompt. By externalizing planning and context management from the LLM, the system keeps context bounded, greatly reduces tool-use failures, and produces explanations that are reproducible and auditable. This separation also addresses production constraints: bounded per-node context scales to large observability datasets, while deterministic control logic minimizes LLM invocations, reducing exposure to API rate limits.
Contributions:
While addressing limitations of existing agent patterns, this paper makes the following contributions:
  • Separation of concerns between deterministic and non-deterministic reasoning: By casting operational diagnosis as structured inference over an explicit operational graph, we shift diagnosis from free-form conversational reasoning tackled by non-deterministic LLMs to the discovery of minimal explanatory subgraphs consistent with observed evidence. This leads to a novel agentic architecture where deterministic and non-deterministic reasoning work in tandem. More specifically, we derive a disaggregated agent architecture in which language models perform bounded local reasoning, while a symbolic controller manages traversal order, context gathering, belief bookkeeping, and termination. This separation enables bounded context, deterministic control flow, reduced tool-use failures, and explicit belief revision.
  • Belief-revision perspective on agentic diagnostics. We identify non-monotonic belief revision as the central challenge in LLM-based diagnostics and show that explanation quality, stability, and ordering are fundamentally coupled through this lens. We connect failures of existing agent patterns to their inability to explicitly represent and revise beliefs under evolving evidence. Drawing inspiration from belief propagation, we introduce a principled agent model in which diagnosis proceeds via iterative local belief assignment and propagation over an operational graph. This model provides a unifying abstraction for context acquisition, reasoning, and explanation in dynamic environments.
  • Validation on operational diagnosis tasks. We validate the proposed formulation and agent model on SRE diagnosis tasks using ITBench Jha et al. (2025), demonstrating improvements in accuracy, stability, and token efficiency over ReAct-style agents. While this paper focuses on operational diagnostics, the same principles have been applied to software bug localization in production systems, in our concurrent work Cui et al. (2025).
@@ 2 Problem Formulation
2.1 The Explanations-over-Graphs (EoG) Problem
We model the environment as an Operational Graph $G=(V,E)$ representing a system topology. In the IT domain, this graph may represent various IT specific components such as applications, services, resources, and code modules. In general, Nodes $V$ represent entities; edges $E$ capture structural relationships (calls, depends-on, contains). Crucially, $E$ encodes topology, not the direction of fault propagation—that must be inferred.
Alerts are events that maps directly to nodes $G$ and we do track them accordingly. Given a set of alerts $A\subseteq V$ (entities exhibiting anomalous behavior), the goal is to discover:
@@ 2.1 The Explanations-over-Graphs (EoG) Problem
    An Explanatory Subgraph $G_{S}=(V_{S},E_{S})$ where $V_{S}\subseteq V$ is the relevant active set and $E_{S}$ are directed edges representing inferred propagation explaining the fault. We use the notation $u\to v$ to mean that “$u$ explains vv”.
@@
    A Frontier $\mathcal{F}\subseteq V_{S}$: the minimal set of origin entities that explain all alerts in $A$.
We use “explanatory” rather than “causal” deliberately: the goal is to infer $G_{S}$ from observational evidence (logs, traces) rather than through interventions. The resulting graph represents the system’s explanation for the observed behavior.
3 Characterizing failure modes of ReAct Agent critical for Operational Diagnostic Tasks
We motivate our approach by characterizing the failure modes of ReAct agents Yao et al. (2023) on ITBench Jha et al. (2025), a benchmark of operational diagnostic tasks. Each ITBench scenario presents an incident in a Kubernetes environment—a service degradation, misconfiguration, or resource failure—and requires the agent to identify the originating cause by querying logs, metrics, traces, and configurations. Scenarios are annotated with ground-truth (GT) entities (the originating cause), non-GT entities (components exhibiting symptoms), and the evidence (log patterns, metric anomalies, error messages) that connects symptoms to their cause. This structure enables post-hoc analysis of where agents fail in the diagnostic pipeline. While we focus on ITBench in this section, the same diagnostic structure—entities, evidence, and explanatory chains—applies to other operational domains, including identifying code defects in production systems Cui et al. (2025) and anomaly explanation in cloud cost management.
We evaluate a range of LLM models (including o4-mini, Gemini 2.5 Pro/Flash, and Claude Opus 4.5) under the ReAct paradigm across multiple runs per task. Our analysis reveals that diagnostic failures stem not merely from lack of domain knowledge, but from structural deficiencies in the ReAct paradigm itself: stochasticity in exploration, brittleness in tool execution, and sensitivity to context ordering.
3.1 The Reliability Gap: Stochasticity vs. Reasoning
@@ 3 Characterizing failure modes of ReAct Agent critical for Operational Diagnostic Tasks
Figure 1: The reliability gap between Pass@$k$ (success in at least one run) and Majority@$k$ (success in the majority of runs). Large gaps indicate stochastic success rather than consistent reasoning.
We first examine absolute diagnostic performance. Across models, mean recall ranges widely: Gemini 3 Flash achieves the highest at 76%, followed by Gemini 3 Pro (50%) and Claude Opus 4.5 (53%). The remaining models fall below 35%, with GPT-OSS-120B at just 12%. These results indicate that, under the ReAct paradigm, most frontier models remain far from deployment-ready on operational diagnostic tasks.
O1:  Mean recall ranges from 12% to 76%. Most models are not deployment-ready.
However, even models with high task completion rates exhibit significant variance across runs. Figure 1 illustrates the gap between Pass@$k$ (success in at least one of $k$ runs) and Majority@$k$ (success in the majority of runs). Gemini 3 Flash achieves 94% Pass@$k$ but drops to 74% Majority@kk—a 20-point gap. Claude Opus 4.5 shows an even larger disparity: 74% Pass@$k$ versus 49% Majority@$k$, with only 34% of scenarios yielding consistent correct diagnoses across all runs. This reliability gap—increasingly recognized as a key challenge in production agentic systems Kartik et al. (2025)—indicates that correct diagnoses often result from favorable exploration paths rather than robust reasoning.
O2:  Large gap between Pass@$k$ and Majority@kk—models succeed by chance, not consistent reasoning.
3.2 Exploration Pathology: The Case for Graph Traversal
(a) GT discovery and assertion failures.
(b) Evidence discovery and use failures.
(c) Exploration vs recall.
Figure 2: Exploration pathology in ReAct agents. (a) Failure to reach the causal entity. (b) Failure to fetch relevant evidence. (c) Correlation between exploration breadth and success. These patterns support modeling diagnosis as iterative graph traversal with bounded context fetching at each hop.
Operational diagnosis is a search problem over a massive state space. Our analysis reveals that ReAct agents fail this search not because they cannot reason, but because they lack the structural priors to navigate system topology.
Origin entity (GT) discovery failures.
The first hurdle is reachability. Figure 2(a) tracks the GT entity through five mutually exclusive stages: GT never saw—the entity never appeared in any tool output (exploration failure); GT saw, not queried—the entity appeared passively but was never actively investigated (attention failure); GT queried, not asserted—the agent investigated but never identified it as the cause (reasoning failure); GT asserted, not in output—the agent identified the GT but omitted it from the final answer (prioritization failure, often due to insufficient evidence or noisy confounding signals); and GT asserted, in output—successful diagnosis. A substantial failure mode is “GT never saw”: in any diagnostic query over a graph—and distributed systems naturally enforce a topology graph—the origin entity is linked to the symptom via dependency edges, yet unstructured data discovery through the ReAct pattern fails to follow these links. Therefore, the diagnostic process must be modeled as a graph traversal, where the agent follows dependency edges from symptoms toward upstream causes.
Evidence discovery failures.
Figure 2(b) shows high incidence of Evidence not in context (red): the necessary logs or metrics existed but were never fetched. This failure mode arises from three distinct causes: (1) the entity containing the evidence was never explored, so the evidence was never seen; (2) tool arguments were too restrictive to retrieve the relevant data (e.g., fetching 5 log lines instead of 50 to conserve context budget); and (3) what constitutes “relevant” data may not be known a priori—an agent cannot know which log lines matter until later evidence reveals their significance (the sequencing problem in §3.4). Evidence ignored (orange) indicates evidence appeared but was never investigated—a reasoning failure. Evidence used (green) indicates successful retrieval and utilization. Notably, the “evidence ignored” category is relatively small for recent frontier models, indicating that modern LLMs understand IT diagnostic semantics and can utilize evidence when present; infusing additional domain knowledge is orthogonal to our contribution. Since agents cannot scan the whole system at once, they must iteratively explore each node’s neighborhood, fetching and analyzing local evidence hop-by-hop.
Structured exploration failures.
Figure 2(c) shows that success rates correlate strongly with exploration breadth: agents exploring fewer than 4 entities achieve below 20% success; those exploring 8+ entities exceed 60%. However, exhaustive exploration is infeasible in production systems with thousands to millions of entities. ReAct agents, which treat every tool call as an independent action, lack mechanisms to systematically expand the frontier of investigated nodes while following evidence. Guided exploration along dependency edges—rather than unstructured data exploration—addresses this gap.
O3:  High rates of “GT never saw” and “Evidence not in context” failures due to lack of structured exploration.
I1:  Enforce structured exploration via iterative graph traversal, guided by evidence.
3.3 Controller Brittleness: The Cost of Interleaved Execution
(a) Plan abandonment rate.
(b) Tool repetition rate.
(c) Tool failure categories.
The ReAct paradigm conflates reasoning (what to investigate) with execution (how to invoke tools). This overloading of the LLM leads to measurable controller instability.
Plan abandonment.
Tool repetition.
Syntactic failures.
Figure 3(c) categorizes tool invocation failures. A substantial fraction are strictly syntactic—malformed arguments, type errors, invalid parameter combinations—rather than semantic reasoning failures Chen et al. (2024b). These errors indicate that LLMs are suboptimal drivers for precise tool orchestration when simultaneously reasoning about complex domains. In contrast, conventional non deterministic programs tend to excel in tool invocation and execution, when they have access to the right contextual information.
O4:  High rates of plan abandonment, tool repetition, and syntactic failures.
I2:  Separation of concerns—decouple Controller (syntax, state) from Reasoner (semantics).
3.4 Path Dependence and Order Sensitivity
Figure 4: Impact of oracle reordering (i.e., presenting relevant entities/evidence earlier in the exploration sequence) on majority@k.
We identify that ReAct agents are highly sensitive to the order in which information is discovered. In an oracle experiment (Figure 4), we reorder the investigation sequence of the agent’s original trajectory to prioritize the path from symptoms to ground-truth entities, without adding any new entities or information, and then re-execute with identical model and tools.
This reordering yields 14–70% improvement in Majority@$k$ across all models, with GPT-OSS-120B showing 70% gains and o4-mini showing 61%. The effect is consistent across model scales and families, indicating a structural property of the paradigm rather than model-specific limitations.
This path dependence confirms that agents struggle with non-monotonic reasoning: they form early beliefs based on initial context and fail to revise them when downstream evidence contradicts upstream assumptions. In diagnostic domains, where the relevance of evidence often becomes apparent only after subsequent discoveries, this rigidity is particularly costly.
O5:  Oracle reordering yields 14–70% improvement in Majority@kk—exploration order materially affects success.
I3:  Belief revision—new evidence must propagate backward to update prior judgments.
3.5 Summary: From Observations to Design Principles
The preceding analysis identifies three structural deficiencies in the ReAct paradigm that limit its efficacy in operational diagnosis:
  1. 1.
    Unguided exploration drives the reliability gap (O2) and exploration failures (O3). The correlation between entity coverage and success motivates I1: use graph topology to guide traversal and ensure systematic coverage.
  2. 2.
    Interleaved reasoning and execution cause controller brittleness (O4). High rates of syntactic failure and plan abandonment motivate I2: delegate tool orchestration and state management to a symbolic controller while restricting the LLM to semantic reasoning over a bounded context.
  3. 3.
    Path-dependent reasoning hinders belief revision (O5). The sensitivity of success rates to investigation order motivates I3: maintain an external state of local assessments that can be revised as new evidence emerges, enabling non-monotonic updates.

### v2

@@ Abstract
LLM agents excel when environments are mostly static, and their required information fits in a model’s context window, but they struggle with diagnostic investigations such as IT incident management, where operators iteratively mine massive observability data to identify the origins that explain observed symptoms, enabling correct remediation. The data in these domains is structured with inherent hidden dependencies: entities interact, signals co-vary, and the importance of a fact may only become clear after other evidence is discovered. To cope with bounded context windows, agents must summarize intermediate findings before their significance is known, increasing the risk of discarding key evidence. ReAct-style/CodeAct agents are especially brittle in this regime. Their retrieve-summarize-reason loop makes conclusions sensitive to exploration order and introduces run-to-run non-determinism, producing a reliability gap where Pass@k may be high, but Majority@k remains low. Increasing samples or reasoning lengths cannot guarantee stable results because, in high-risk diagnostic scenarios, consistent accuracy is required over sporadic success. Furthermore, the ReAct framework lacks a mechanism to revise its beliefs as new information becomes available. In addition, ReAct entangles semantic reasoning with controller duties such as tool orchestration and state tracking; execution errors and plan drift degrade reasoning while consuming scarce context. We address these issues by formulating the investigation as abductive reasoning over a dependency graph and proposing EoG (Explanations over Graphs), an agentic architecture where an LLM performs bounded local evidence mining and labeling (cause vs symptom) while a deterministic controller manages traversal, state, and belief propagation to compute a minimal explanatory frontier. On representative ITBench diagnostics tasks, EoG improves accuracy and run-to-run consistency over ReAct baselines, including a 7×  average gain in Majority@k F1 score.
Preprint. Work in progress.
1 Introduction
LLM agents excel when environments are mostly static and their required information fits in the context window. They struggle with diagnostic investigations such as incident management, where operators iteratively mine massive observability data to identify the origins that explain observed symptoms, enabling correct remediation. This challenge affects numerous fields, including medical diagnosis Varatharajah et al. (2017) and security threat analysis Wynn (2014). It is particularly critical in incident management, where cloud infrastructure and microservices constantly generate massive amounts of logs, metrics, traces, and events. When failures occur, operators observe symptoms (alerts, errors, degradations) but must identify the minimal set of upstream sources whose effects propagate through the environment to explain observations; only then can they remediate. Since remediation based on incorrect diagnosis carries significant operational risk, the goal is to infer explanations from observational evidence rather than through intervention. More formally, this is abductive reasoning over massive observability data from an environment with inherent graphical structure composed of entities connected by dependency edges (§2).
Limits of existing agent patterns: Current agentic approaches struggle with abductive reasoning. For example, on the SRE scenarios of ITBench Jha et al. (2025) [^1] ITBench is an open-source benchmark for testing how well AI agents can diagnose and resolve IT issues. (pertaining to incident management), the best ReAct agent achieves only 13.81% pass@1 recall (§3). Beyond well-known limitations such as tool invocation failures and redundant or degenerate tool calls  Cemri et al. (2025), ReAct Yao et al. (2023) and CodeAct Wang et al. (2024)-based agents struggle significantly with context gathering and action sequencing under uncertainty. Reasoning is only as sound as the evidence it conditions on: if relevant context is not retrieved, it cannot be reasoned over. In open-ended investigations, evidence relevance is not known a priori, and ReAct agents lack mechanisms for systematic context acquisition. Tool calls are issued opportunistically, missing low-frequency, high-impact signals or retrieving excessive irrelevant data. Summarizing the retrieved evidence before reasoning introduces a critical failure mode that discards information before relevance can be established, often losing critical evidence  Liu et al. (2024).
Furthermore, open-ended investigations exhibit non-monotonic belief revision behaviors: interpretation of entities with the related signals naturally changes as new evidence emerges elsewhere in the system, making reasoning order-dependent. The sequence of evidence retrieval, summarization, and presentation to the LLM materially affects conclusions. In practice, ReAct agents show high variance across runs due to planning non-determinism, unstable exploration orderings, and lossy summarization, producing fragile, incomplete, irreproducible explanations. These failures are architectural: ReAct implicitly assumes linear, monotonic reasoning over an accumulating context, whereas abductive reasoning over massive graphs requires structured context acquisition with explicit belief revision.
Towards graph-guided agentic explanations: We introduce an operational-graph abstraction for investigations under partial observability. We represent the environment as an operational graph and define diagnosis as explanations over such dynamically evolving operational graphs: explanations correspond to minimal subgraphs consistent with observed evidence. Framing diagnosis in this way converts an ill-defined exploration problem into structured inference. For example, across IT domains, abductive explanations are naturally structured by dependency relationships among system entities.
"Diagnosis is non-monotonic because new evidence can change what we believe about an entity’s role. Drawing from belief propagation Pearl (1988), we model diagnosis as an iterative local belief assignment, revision, and propagation process through the graph. This leads to a disaggregated architecture: local reasoning occurs via LLMs over bounded neighborhoods, while a deterministic controller manages traversal order, context gathering, belief bookkeeping, and termination. This separation enables non-monotonic belief revision through explicit recording and updating of beliefs rather than implicit overwriting in growing ReAct-style prompts. By externalizing planning and context management, the system keeps context bounded, reduces tool-use failures, and produces reproducible, auditable explanations. This separation also addresses production constraints: bounded per-node context scales to large datasets, while deterministic control minimizes LLM calls and API rate limit exposure. Our contributions:
  • EoG agent with separation of concerns: We develop EoG (Explanations over Graphs), a disaggregated architecture where a Deterministic Controller manages graph traversal, state, and tool orchestration, while a stateless Abductive Policy $\pi_{\text{abd}}$ classifies each node as origin, symptom, or healthy based on bounded local evidence (§4). This separation eliminates controller failures, ensures auditability, and enables systematic exploration.
  • Semantic Belief Propagation: We introduce SBP, a message-passing mechanism inspired by belief propagation Pearl (1988), to address non-monotonic belief revision. When new evidence changes a node’s belief, it broadcasts to neighbors, triggering reconsideration by $\pi_{\text{abd}}$—correcting early misdiagnoses without restart (§4.3).
@@ 2 Problem Formulation
We model the environment as a Dependency Graph $G=(V,E)$ where nodes $V$ are entities and edges $E$ encode structural dependencies along which effects propagate. Entity examples include IT services in SRE settings, hosts in security, or physiological signals in medical diagnosis. Entities can originate faults, exhibit symptoms, or propagate effects along domain dependency edges. In this work, we focus on abductive reasoning in SRE to diagnose failures from observed anomalies, where $V$ are services, pods, and configurations, and $E$ are API calls or ownership relations.
An anomaly is an observation of abnormal behavior at an entity; in SRE, alerts are anomalies that operators configure to receive notifications. Given anomalies $\mathcal{O}$ on entities $V_{\mathcal{O}}\subseteq V$, the task is to explain them. Formally, we seek:
@@
    An Explanatory Subgraph $G_{S}=(V_{S},E_{S})$: $V_{S}\subseteq V$ is the active set; $E_{S}$ are directed edges representing inferred propagation: $u\to v$ means “$u$ explains vv”.
@@
    A Frontier $\mathcal{F}\subseteq V_{S}$: the minimal origin set whose faults explain all anomalies on $V_{\mathcal{O}}$.
Our focus is explanatory inference from observational evidence. We avoid interventions since remediation actions carry operational risk. This formulation naturally extends to other domains such as FinOps, DevOps, healthcare, and forensics.
3 Characterizing Failure Modes
We motivate our approach by characterizing failure modes of ReAct agents Yao et al. (2023) built using Codex Wang et al. (2024); OpenAI (2025) (Appendix A) on the SRE scenarios of ITBench Jha et al. (2025).We focus on SRE scenarios as they require abductive reasoning—inferring root causes from observed symptoms.
3.1 The Reliability Gap
@@
Figure 1: The reliability gap between Pass@$k$ (success in at least one run) and Majority@$k$ (success in the majority of runs). Significant deviations from dotted line indicate stochastic success rather than consistent reasoning.
While most models perform poorly under the ReAct paradigm, we quantify a more fundamental issue: stochastic success rather than consistent reasoning. We define the reliability gap as Pass@$k$ (success in at least one of $k$ runs) minus Majority@$k$ (success in the majority of runs). [^2] When verification is cheap (e.g., code generation with test suites), one can sample until a correct answer is verified, making Pass@$k$ sufficient for accuracy. However, even then, low Majority@$k$ has consequences for maintainability, latency, and predictability. In incident management, verification requires remediation actions that carry operational risk, so consistent correctness is essential. Figure 1 shows this gap ranging from 14 points for Gemini 3 Flash to 32 points for Gemini 2.5 Pro. Even the best model (Gemini 3 Flash: 89% Pass@$k$, 74% Majority@$k$) exhibits significant variance, with only 51% of scenarios yielding consistent correct diagnoses across all runs. This reliability gap must be addressed to ensure successful deployment of agents in enterprise settings Kartik et al. (2025).
3.2 Exploration Failures
(a) GT discovery funnel.
(b) Evidence discovery failures.
(c) Oracle reordering impact.
Figure 2: ReAct agent failure modes. (a) GT discovery funnel: never saw (exploration), saw not queried (attention), queried not asserted (reasoning), asserted not in output (prioritization), asserted in output (success). (b) Evidence discovery: not in context (red), ignored (orange), used (green). (c) Impact of oracle reordering on Majority@$k$.
Abductive reasoning in complex systems is reduced to a search over a dependency graph of entities. Given observed symptoms, the goal is to identify the minimal upstream entities whose states explain all observations. This search space is exponential; the dependency graph constrains tractable exploration. Our analysis reveals that ReAct agents fail not because they cannot reason, but because they lack structural priors to use a graph.
Ground truth (GT) entity discovery failures. Figure 2(a) tracks the GT entity through a discovery funnel. A substantial failure mode is “GT never saw”: the GT entity is linked to the symptom via dependency edges, yet the ReAct agent fails to follow these links.
Evidence discovery failures. Figure 2(b) shows a high incidence of Evidence not in context: supporting data existed but was never fetched, either because the entity was never explored or because what constitutes “relevant” evidence depends on information discovered later (§2). The small “evidence ignored” category indicates that modern LLMs can use evidence when it is present; the bottleneck is discovery, not reasoning.
Path dependence. ReAct agents are highly sensitive to information ordering. In an oracle experiment (Figure 2(c)), reordering investigation sequences to prioritize paths from symptoms to root causes without adding new information, yields 14–70% improvement in Majority@$k$ across LLMs. This path dependence indicates agents struggle with non-monotonic reasoning: early wrong beliefs persist despite emerging evidence.
3.3 Controller Failures
ReAct agents experience frequent controller failures—plan abandonment, tool repetition, and syntactic errors. We use MAST UC Berkeley (2025) to characterize these failures on our agent trajectories (Appendix C). The key implication is the separation of deterministic control flow (tool orchestration, state management) from non-deterministic reasoning (causal inference).